# Disentangling spatio-temporal knowledge for weakly supervised object detection and segmentation in surgical video

### v1

@@ 4 Experiments
4.1 Datasets
To investigate the characteristics of our proposed method, we utilize two distinct video dataset collections: YouTube-Objects video dataset (v2.2) [17] and Cholec80 dataset [43]. YouTube-objects data contains 106 videos for training, and 49 videos for testing, where ten categories ($N=10$) of objects are covered. Each video is split into short-duration clips, and in each clip a few frames are annotated with a bounding box for evaluation. The Cholec80 dataset consists of 80 cholecystectomy surgery videos, each downsampled to 1 FPS, with labels indicating the presence or absence of seven surgical tools ($N=7$) in each frame. We create a variant TOP video dataset based on Cholec80. First these videos are split into 30-second clips without overlap, merging the tool labels across frames to generate a video-level tool presence binary vector label for each clip. The test set is based on CholecSeg8K [11] which contains 8K images from 17 videos in Cholec80 with segmentation labels for instruments and anatomy. From this dataset we derive 100 video clips containing 3K frames. To prevent contamination between testing and training data, 5,296 training clips are sampled from from 63 Cholec80 videos not included in CholecSeg8k. We calculate the tool frequencies across clips using frame per clip (FPC) which denotes the percentage of frames in a clip with the tool present. YouTube-Objects dataset has nearly 100% FPC for each clip and thus represents the second case (Type II – WSVOS-COP) demonstrated in Figure 1. Most categories in Cholec80 based dataset appear on average in less than 60% of the frames of a particular video labeled with that category, which makes it a much more challenging dataset that represents a WSVOS-TOP (Type III) problem.
4.2 Experimental Setup and Evaluation Metrics
@@
4.3 Analysis on Youtube video object detection Perfomance
Our method is primarily designed for ViT-based feature extractors, utilizing the pre-trained DINO ViT [2] as our backbone and fine-tuning are on domain-specific data. We also measure performance using a ResNet34 backbone with our proposed spatio-temporal disentanglement methodology, by converting ResNet features to match the format produced by ViT. We investigate different optimization strategies for the backbone on new domain data: back-propagation through the teacher (denoted as “refined by T”), through the student (“refined by S”), or keeping the backbone frozen. Detection outputs from either Teacher (T) or final student (full) or extracted from fused CAM (T-S fusion) are presented. Our experiments show that even a completely frozen ResNet captures a certain level of generalized features. Through our semi-decoupled KD approach, we achieve a performance improvement from 56.6% to 69.1%. Interestingly, refining the backbone through the student tends to degrade the final output compared to MLP teacher (68.4% to 67.4%), likely due to the challenge of learning generalized features from video-level labels with higher uncertainty. Optimizing the backbone through the teacher proves to be the effective approach for both ResNet and ViT backbones, leading to a stronger student model by leveraging well-learned features. For ResNet, we observe that the T-S fusion has better bounding box localization than final student (72.2% vs. 70.6%), possibly because ResNet features are less generalized and lean more toward the MLP from the teacher module, while the activations learned by the student complement those of the teacher. In conclusion, we recommend using ViT within our framework and advise against direct student-driven backbone refinements.
4.4 Surgical video tool segmentation performance
Table 2: Comparison to state-of-the-art weakly supervised methods with either frame-level supervision or video-level supervision. Video (V) and Frame (F) accuracy (AC)
MCTformer [45] 98.27 97.63 43.14 50.07
LayerCAM[15] 95.67 94.87 31.97 75.59
XGradCAM[5] 92.21 90.21 32.13 53.53
TCAM [1] 97.08 95.10 28.22 54.60
T only (ours) 98.26 97.56 47.50 48.72
T-S fusion (ours) 97.11 97.23 48.64 39.61
VDST-Net (ours) 96.82 98.23 61.80 28.10
@@ 4.4 Surgical video tool segmentation performance
On surgical videos, we compare our method to frame-level weakly supervised methods MCTformer [45], LayerCAM [15], and XGradCAM [5], all trained with frame-level labels. Models were adapted to our dataset using their open-source code implementations. We also compare our method to the video-level supervised method TCAM [1], where both our method and TCAM are trained with more challenging video-level class labels. We test our model using teacher (T) only, teacher-student (T-S) fusion, and full VDST-Net variations. For the surgical video data we utilize another pre-trained Masked Auto-Encoder (MAE) ViT backbone [9] which is tuned on a large scale data for segmentation[18] before it is adapted to the new domain (Cholec80) data.
4.5 Ablation Study
@@
Effect of ranked top-$k$ pooling:   We investigated the effects of transitioning from ranked top-$k$ pooling to average and max pooling, observing a decline in segmentation performance. Specifically, with average pooling, we noted a decrease in Dice score by 11.49%, and with max pooling, the decrease was 11.71%. Furthermore, there was a considerable decrease in frame classification accuracy, amounting to 3.72% with average pooling and 3.04% with max pooling.
TPC decoupling strategy:   Additionally, the proposed TPC slice kernel for selecting outsignificant activations, enhances the student’s learning. The absence of this component leads to a decline in segmentation performance by 4.16% in Dice and 1.34% in IoU, as demonstrated in Table 3, due to over-introduced loss from non-target class knowledge [52].
@@ 5 Conclusion
This work was supported by the University of Pennsylvania Thomas B. McCabe and Jeannette E. Laws McCabe Fellow Award and the Linda Pechenik Montague Investigator Award

### v3

@@ 4.4 Analysis on Youtube video object detection Perfomance
We utilize the pre-trained DINO ViT[2] as backbone and fine-tune it on domain-specific data. We also measure performance using a ResNet34 backbone by converting features to the form of ViT. We investigate different optimization strategies for the backbone on new domain data: back-propagation through the teacher (denoted as “refined by T”), through the student (“refined by S”), or keeping the backbone frozen. Detection outputs from either Teacher (T) or final student (full) or extracted from fused CAM (T-S fusion) are presented. Our experiments show that even a frozen ResNet captures a certain level of generalized features. Through our semi-decoupled KD approach, we improve performance from 56.6% to 69.1%. Interestingly, refining the backbone through the student tends to degrade the final output compared to MLP teacher (68.4% to 67.4%), likely due to the challenge of learning generalized features from video-level labels. Optimizing the backbone through the teacher proves to be the effective approach for both ResNet and ViT backbones leveraging well-learned features. For ResNet, we observe that the T-S fusion has a better localization score than the final student (72.2% vs. 70.6%), possibly because ResNet features are less generalized and lean more toward the MLP from the teacher module, while the activations learned by the student complement those of the teacher. In conclusion, we recommend using ViT within our framework and advise against direct student-driven backbone refinements.
4.5 Ablation Study
@@
Effect of ranked top-$k$ pooling:   We compare the ranked top-$k$ pooling to global average and max pooling. Specifically, with average pooling, we noted a decrease in Dice score by 11.49%, and with max pooling, the decrease was 11.71%. Furthermore, there was a considerable decrease in frame classification accuracy, amounting to 3.72% with average pooling and 3.04% with max pooling.
TPC decoupling strategy:   Additionally, the proposed TPC slice kernel for activations gating, enhances the student’s learning. The absence of this component leads to a decline in segmentation performance by 4.16% in Dice and 1.34% in IoU, as demonstrated in Table 3, due to over-introduced loss from non-target class knowledge [57].
@@ 5 Conclusion
Anonymized Acknowledgements.